# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

### v1

@@ Abstract
Bi-factor analysis is a form of confirmatory factor analysis widely used in psychological and educational measurement. The use of a bi-factor model requires the specification of an explicit bi-factor structure on the relationship between the observed variables and the group factors. In practice, the bi-factor structure is sometimes unknown, in which case an exploratory form of bi-factor analysis is needed to find the bi-factor structure. Unfortunately, there are few methods for exploratory bi-factor analysis, with the exception of a rotation-based method proposed in Jennrich and Bentler, (2011, 2012). However, this method only finds approximate bi-factor structures, as it does not yield an exact bi-factor loading structure, even after applying hard thresholding. In this paper, we propose a constraint-based optimisation method that learns an exact bi-factor loading structure from data, overcoming the issue with the rotation-based method. The key to the proposed method is a mathematical characterisation of the bi-factor loading structure as a set of equality constraints, which allows us to formulate the exploratory bi-factor analysis problem as a constrained optimisation problem in a continuous domain and solve the optimisation problem with an augmented Lagrangian method. The power of the proposed method is shown via simulation studies and a real data example. Extending the proposed method to exploratory hierarchical factor analysis is also discussed. The codes are available on “https://anonymous.4open.science/r/Bifactor-ALM-C1E6”.
@@ 1 Introduction
The bi-factor model was originally proposed by Holzinger and Swineford, (1937) for linear factor analysis and further extended by Gibbons and Hedeker, (1992), Gibbons et al., (2007), Cai et al., (2011), among others, to nonlinear factor analysis settings to account for dichotomous, ordinal, and nominal data. These models assume the observed variables can be accounted for by $(G+1)$ factors, with a general factor, onto which all items load directly, and $G$ group factors that each is associated with a subset of variables. Such a specification leads to good interpretations in many real-world applications. These models have received wide applications in psychological and educational measurement; see (e.g., Bradlow et al.,, 1999; Cai et al.,, 2016; Chen et al.,, 2012; DeMars,, 2006, 2012; Gibbons et al.,, 2009; Gignac and Watkins,, 2013; Jeon et al.,, 2013; Reise et al.,, 2007; Rijmen,, 2010). However, we note that all these applications of bi-factor analysis are confirmatory, in the sense that one needs to pre-specify the number of group factors and the relationship between the observed variables and the group factors. Such prior knowledge may not always be available. In that case, an exploratory form of bi-factor analysis is needed.
@@
However, standard exploratory factor analysis methods are not applicable to the bi-factor analysis setting, and few methods have been developed for exploratory bi-factor analysis. An exception is the seminal work of Jennrich and Bentler, (2011, 2012), who proposed a rotation-based method for exploratory bi-factor analysis with orthogonal and oblique factors, respectively. However, their approach has some limitations. First, as a common issue with rotation-based methods, their method does not yield many zero loadings, and thus, the resulting loading structure does not have an exact bi-factor structure. Although a post-hoc thresholding procedure (i.e., treating loadings with an absolute value below a threshold as zero) can be applied to obtain a cleaner loading pattern, it does not work well when some variables show relatively large loadings on more than one group factor after the rotation. In fact, one cannot always find a threshold that yields an exact bi-factor structure that each variable loads on one and only one group factor. Second, as noted in Jennrich and Bentler, (2012), their method fails completely in the best case where there is a rotation of an initial loading matrix that has an exact bi-factor structure. This failure is due to that their rotation method cannot incorporate the zero constraints on the correlations between the general factor and the group factors.
@@
This paper proposes a constrained optimisation method for exploratory bi-factor analysis, which overcomes the issues with the rotation-based method. The contribution is four-fold. First, we provide a mathematical characterisation of the bi-factor loading structure as a set of nonlinear equality constraints, which allows us to formulate the exploratory bi-factor analysis problem as a constrained optimisation problem. In other words, it turns a discrete model selection problem into a continuous optimisation problem, which reduces the computational demand in some sense. It is shown that in the aforementioned best case where the rotation method fails, the global solutions to the optimisation can perfectly recover the exact bi-factor structure. Second, we propose an augmented Lagrangian method (ALM, Bertsekas,, 2014) for solving this optimisation problem, which is a standard numerical optimisation method for solving constrained optimisation with robust empirical performance and good theoretical properties. Third, we combine the proposed method with the Bayesian information criterion (BIC, Schwarz,, 1978) for selecting the number of group factors. Compared with existing exploratory factor analysis methods for determining the number of factors, our method is tailored to the bi-factor model structure and, thus, tends to be statistically more efficient when the data is indeed generated by a bi-factor model. Finally, we demonstrate that the proposed method can be extended to learning the loading structure of hierarchical factor models (Schmid and Leiman,, 1957; Yung et al.,, 1999), a higher-order extension of the bi-factor model that has received wide applications (see, e.g., Brunner et al.,, 2012, and references therein). The bi-factor model can be viewed as a special hierarchical factor model with a two-layer factor structure, with the general factor in one layer and the group factors in the other. Similar to exploratory bi-factor analysis, the proposed method yields exact hierarchical factor loading structures without a need for post-hoc treatments.
@@
The rest of the paper is organised as follows. In Section 2, we formulate the exploratory bi-factor analysis problem as a constrained optimisation problem and propose an ALM for solving it. We also propose a BIC-based procedure for selecting the number of group factors. Simulation studies and a real data example are presented in Sections 3 and 4, respectively, to evaluate the performance of the proposed method. We conclude with discussions in Section 5. An extension of the proposed method to exploratory hierarchical factor analysis is given in the Appendix, including the extended method and a simulation study.
@@ 2.1 Bi-factor Model Structure and a Constrained Optimisation Formulation
For the ease of exposition and simplification of the notation, we focus on the linear bi-factor model, while noting that the constraints that we derive for the bi-factor loading matrix below can be combined with the likelihood function of other bi-factor models (e.g., Gibbons and Hedeker,, 1992; Gibbons et al.,, 2007; Cai et al.,, 2011) for their exploratory analysis. We focus on the extended bi-factor model, also known as the oblique bi-factor model, as considered in Jennrich and Bentler, (2012) and Fang et al., (2021). This model is more general than the standard bi-factor model, in the sense that the latter assumes all the factors to be uncorrelated while the former allows the group factors to be correlated. As established in Fang et al., (2021), this extended bi-factor model is identifiable under mild conditions. We note that the proposed method can be easily adapted to the standard bi-factor model.
Consider a dataset with $N$ observation units from a certain population and $J$ observed variables. The extended bi-factor model assumes that the $J$ variables are partitioned into $G$ non-overlapping groups. Let $\mathcal{B}_{g}\subset\{1,...,J\}$ denote the set of variables in the $g$th group. Then the bi-factor model assumption implies that $\mathcal{B}_{g}\cap\mathcal{B}_{g^{\prime}}=\emptyset$, $g\neq g^{\prime}$, and $\cup_{g=1}^{G}\mathcal{B}_{g}=\{1,...,J\}$. The model further assumes that the population covariance matrix of the observed variables can be decomposed as
@@
where $\Lambda=(\lambda_{jg})_{J\times(G+1)}$ is the loading matrix, $\Phi=(\phi_{gg^{\prime}})_{(G+1)\times(G+1)}$ is the correlation matrix of the factors, which is assumed to be strictly positive definite, and $\Psi$ is a $J\times J$ diagonal matrix with diagonal entries $\psi_{1}$, …, $\psi_{J}$. The bi-factor structure imposes the following zero constraints on the loading matrix $\Lambda$:
@@
Now suppose that the number of group factors $G$ is known, while the variable partition, $\mathcal{B}_{g},g=1,...,G$, is unknown. Section 2.3 considers the selection of $G$ when it is unknown. The bi-factor structure means that for each $j$, there is at most one non-zero element in $(\lambda_{j,2},...,\lambda_{j,G+1})^{\top}$. Consequently, the loading matrix $\Lambda$ should satisfy the following $J(G-1)G/2$ constraints:
$\lambda_{jk}\lambda_{jk^{\prime}}=0,\mbox{~{}for all~{}}k,k^{\prime}=2,...,G+1,k\neq k^{\prime},j=1,...,J$ (1)
@@
where $l$ is a loss function and $S$ is the sample covariance matrix of observed data. We focus on the case when $l$ is the fit function based on the normal likelihood (Chen et al.,, 2023)
@@
while noting that this loss function can be replaced by other loss functions for factor analysis (see, e.g., Chen et al.,, 2023), including the Frobeneious norm of $\Lambda\Phi\Lambda^{\top}+\Psi-S$ that is used in the least square estimator for factor analysis.
The following theorem shows that the proposed method can perfectly recover the true bi-factor loading structure in the best case when $S=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$ is the true covariance matrix of data, where $\Lambda^{*}$, $\Phi^{*}$ and $\Psi^{*}$ denote the true value of the corresponding parameter matrices. Note that the rotation method proposed in Jennrich and Bentler, (2012) completely fails in this case.
@@
The following conditions are sufficient for the identifiability of the bi-factor structure and its parameters.
@@
Given $S=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$. Suppose that there exists another pair of parameters $\Lambda,\Phi,\Psi$ satisfy the bi-factor structure constraints, we have $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}=\Lambda\Phi(\Lambda)^{\top}$ and $\Psi^{*}=\Psi$.
@@
$|\mathcal{H}^{*}|\geq 2$. In addition, there exists $g_{1}\in\mathcal{H}^{*}$ such that $|Q_{g_{1}}^{*}|\geq 3$ and any 2 rows of $\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}]$ are linearly independent.
@@
Suppose that Conditions 1-2 hold. For any parameters $\Lambda,\Phi,\Psi$ that satisfy $S=\Lambda\Phi(\Lambda)^{\top}+\Psi$, there exist a diagonal sign-flip matrix $D\in\mathcal{D}$ and a permutation matrix $P\in\mathcal{P}$ such that $\Lambda=\Lambda^{*}PD$ and $\Phi=DP^{\top}\Phi^{*}PD$.
The Proof of Theorem 1 is given in the appendix.
@@
We note that without additional information, the best we can achieve is to recover $\Lambda^{*}$ and $\Psi^{*}$ up to $\Lambda=\Lambda^{*}PD$ and $\Psi=DP^{\top}\Psi^{*}PD$, where the permutation matrix $P$ and sign-flip matrix $D$ are necessary to account for the label and sign indeterminacies of the factor model. In that case, $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}=\Lambda\Phi(\Lambda)^{\top}$.
@@
Condition 1 ensures the separation between low rank matrix $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$ and diagonal matrix $\Psi^{*}$. One of the sufficient condition of Condition 1 that can be easily checked in practice is given in Condition 3 below. We refer to Theorem 5.1 in Anderson and Rubin, (1956) and Theorem 2 in Fang et al., (2021) for further choices of conditions to replace Condition 1.
@@
$|\mathcal{Q}^{*}_{g}|\geq 3$ for all $g=1,\ldots,G$ and $|\mathcal{H}^{*}|\geq 3$.
@@
Conditions 2 is slightly different with the Condition E3S for Proposition 1 in Fang et al., (2021), which shows the identifiability of parameters when the bi-factor structure is known. Here, combined with Condition 1, we show that such conditions are sufficient to ensure the identifiability under an unknown bi-factor structure.

### v2

@@ Abstract
Bi-factor analysis is a form of confirmatory factor analysis widely used in psychological and educational measurement. The use of a bi-factor model requires the specification of an explicit bi-factor structure on the relationship between the observed variables and the group factors. In practice, the bi-factor structure is sometimes unknown, in which case an exploratory form of bi-factor analysis is needed to find the bi-factor structure. Unfortunately, there are few methods for exploratory bi-factor analysis, with the exception of a rotation-based method proposed in Jennrich and Bentler, (2011, 2012). However, the rotation-based method only finds approximate bi-factor structures, as it does not yield an exact bi-factor loading structure, even after applying hard thresholding. In this paper, we propose a constraint-based optimisation method that learns an exact bi-factor loading structure from data, overcoming the issue with the rotation-based method. The key to the proposed method is a mathematical characterisation of the bi-factor loading structure as a set of equality constraints, which allows us to formulate the exploratory bi-factor analysis problem as a constrained optimisation problem in a continuous domain and solve the optimisation problem with an augmented Lagrangian method. The power of the proposed method is shown via simulation studies and a real data example. Extending the proposed method to exploratory hierarchical factor analysis is also discussed. Code implementing the proposed method is open-source and publicly available at https://anonymous.4open.science/r/Bifactor-ALM-method-757D.
@@ 1 Introduction
The bi-factor model was originally proposed by Holzinger and Swineford, (1937) for linear factor analysis and further extended by Gibbons and Hedeker, (1992), Gibbons et al., (2007), Cai et al., (2011), among others, to nonlinear factor analysis settings to account for dichotomous, ordinal, and nominal data. These models assume that the observed variables can be accounted for by $(G+1)$ factors, with a general factor, onto which all items load directly, and $G$ group factors that each is associated with a subset of variables. Such a specification leads to good interpretations in many real-world applications. These models have received wide applications in psychological and educational measurement; see e.g., Bradlow et al., (1999); Cai et al., (2016); Chen et al., (2012); DeMars, (2006, 2012); Gibbons et al., (2009); Gignac and Watkins, (2013); Jeon et al., (2013); Reise et al., (2007); Rijmen, (2010). However, we note that all these applications of bi-factor analysis are confirmatory in the sense that one needs to pre-specify the number of group factors and the relationship between the observed variables and the group factors. Such prior knowledge may not always be available. In that case, an exploratory form of bi-factor analysis is needed.
Exploratory bi-factor analysis can be seen as a special case of exploratory factor analysis, which dates back to the seminal work of Thurstone, (1947) concerning finding a “simple structure” of loadings. Various rotation methods have been proposed for exploratory factor analysis. A short list of relevant works includes Kaiser, (1958), McCammon, (1966), Jennrich and Sampson, (1966), McKeon, (1968), Crawford and Ferguson, (1970), Yates, (1987), Jennrich, (2006), Jennrich, (2004), Jennrich and Bentler, (2011, 2012), and Liu et al., (2023). We refer the readers to Browne, (2001) for a review of rotation methods for exploratory factor analysis.
@@
However, standard exploratory factor analysis methods do not apply to the bi-factor analysis setting, and few methods have been developed for exploratory bi-factor analysis. An exception is the seminal work of Jennrich and Bentler, (2011, 2012), who proposed a rotation-based method for exploratory bi-factor analysis with orthogonal or oblique factors. However, their approach has some limitations. First, as a common issue with rotation-based methods, their method does not yield many zero loadings, and thus, the resulting loading structure does not have an exact bi-factor structure. Although a post-hoc thresholding procedure (i.e., treating loadings with an absolute value below a threshold as zero) can be applied to obtain a cleaner loading pattern, it does not work well when some variables show relatively large loadings on more than one group factor after the rotation. In fact, one cannot always find a threshold that yields an exact bi-factor structure that each variable loads on one and only one group factor. Second, as noted in Jennrich and Bentler, (2012), their method fails completely in the best case where there is a rotation of an initial loading matrix that has an exact bi-factor structure. This failure is due to that their rotation method cannot incorporate the zero constraints on the correlations between the general factor and the group factors.
@@
This paper proposes a constrained optimisation method for exploratory bi-factor analysis, which overcomes the issues with the rotation-based method. The contribution is four-fold. First, we provide a mathematical characterisation of the bi-factor loading structure as a set of nonlinear equality constraints, which allows us to formulate the exploratory bi-factor analysis problem as a constrained optimisation problem. In other words, it turns a discrete model selection problem into a continuous optimisation problem, which reduces the computational demand in some sense. It is shown that in the aforementioned best case where the rotation method fails, the global solutions to the optimisation can perfectly recover the true bi-factor structure. Second, we propose an augmented Lagrangian method (ALM, Bertsekas,, 2014) for solving this optimisation problem, which is a standard numerical optimisation method for solving constrained optimisation with robust empirical performance and good theoretical properties. Third, we combine the proposed method with the Bayesian information criterion (BIC, Schwarz,, 1978) for selecting the number of group factors. Compared with existing exploratory factor analysis methods for determining the number of factors, our method is tailored to the bi-factor model structure and, thus, tends to be statistically more efficient when the data is indeed generated by a bi-factor model. Finally, we demonstrate that the proposed method can be extended to learning the loading structure of hierarchical factor models (Schmid and Leiman,, 1957; Yung et al.,, 1999), a higher-order extension of the bi-factor model that has received wide applications (see, e.g., Brunner et al.,, 2012, and references therein). The bi-factor model can be viewed as a special hierarchical factor model with a two-layer factor structure, with the general factor in one layer and the group factors in the other. Similar to exploratory bi-factor analysis, the proposed method yields exact hierarchical factor loading structures without a need for post-hoc treatments.
@@
The rest of the paper is organised as follows. In Section 2, we formulate the exploratory bi-factor analysis problem as a constrained optimisation problem and propose an ALM for solving it. We also propose a BIC-based procedure for selecting the number of group factors. Simulation studies and a real data example are presented in Sections 3 and 4, respectively, to evaluate the performance of the proposed method. We conclude with discussions in Section 5. The appendix in the supplementary material includes additional details about the computation, the simulation studies and the real data example, an extension of the proposed method to exploratory hierarchical factor analysis, and proof of the theoretical results.
@@ 2.1 Bi-factor Model Structure and a Constrained Optimisation Formulation
For the ease of exposition and simplification of the notation, we focus on the linear bi-factor model, while noting that the constraints that we derive for the bi-factor loading matrix below can be combined with the likelihood function of other bi-factor models (e.g., Gibbons and Hedeker,, 1992; Gibbons et al.,, 2007; Cai et al.,, 2011) for their exploratory analysis. We focus on the extended bi-factor model, also known as the oblique bi-factor model, as considered in Jennrich and Bentler, (2012) and Fang et al., (2021). This model is more general than the standard bi-factor model, in the sense that the latter assumes all the factors to be uncorrelated while the former allows the group factors to be correlated. As established in Fang et al., (2021), this extended bi-factor model is identifiable under mild conditions. We also point out that the proposed method can be easily adapted to the standard bi-factor model.
Consider a dataset with $N$ observation units from a certain population and $J$ observed variables. The extended bi-factor model assumes that there exists a general factor and $G$ group factors. The group factors are loaded by independent clusters of variables, where each variable belongs to only one cluster. The model further assumes that the population covariance matrix of the observed variables can be decomposed as
@@
where $\Lambda=(\lambda_{jg})_{J\times(G+1)}$ is the loading matrix, $\Phi=(\phi_{gg^{\prime}})_{(G+1)\times(G+1)}$ is the correlation matrix of the factors, which is assumed to be strictly positive definite, and $\Psi$ is a $J\times J$ diagonal matrix with diagonal entries $\psi_{1}$, …, $\psi_{J}$. Let $\mathcal{B}_{g}\subset\{1,...,J\}$ denote the cluster of variables loading on the $g$th group factor. Then the bi-factor model assumption implies that $\mathcal{B}_{g}\cap\mathcal{B}_{g^{\prime}}=\emptyset$, $g\neq g^{\prime}$, and $\cup_{g=1}^{G}\mathcal{B}_{g}=\{1,...,J\}$. The following zero constraints on the loading matrix $\Lambda$ hold:
@@
Now suppose that the number of group factors $G$ is known, while the clusters $\mathcal{B}_{g},g=1,...,G$, are unknown. Section 2.3 considers the selection of $G$ when it is unknown. The bi-factor structure means that for each $j$, there is at most one non-zero element in $(\lambda_{j,2},...,\lambda_{j,G+1})^{\top}$. Consequently, the loading matrix $\Lambda$ should satisfy the following $J(G-1)G/2$ constraints:
$\lambda_{jk}\lambda_{jk^{\prime}}=0,\mbox{~{}for all~{}}k,k^{\prime}=2,...,G+1,k\neq k^{\prime},j=1,...,J.$ (1)
@@
where $l$ is a loss function and $S$ is the sample covariance matrix of observed data. We focus on the case when $l$ is the fit function based on the normal likelihood
@@
while noting that this loss function can be replaced by other loss functions for factor analysis (see, e.g., Chen et al.,, 2023), including the Frobeneious norm of $\Lambda\Phi\Lambda^{\top}+\Psi-S$ that is used in the least square estimator for factor analysis. We can also replace the sample covariance matrix in (2) with the sample correlation matrix, which is equivalent to performing exploratory bi-factor analysis based on variance-standardised variables.
The following theorem shows that the proposed method can perfectly recover the true bi-factor loading structure in the best case when $S=\Sigma^{*}$, where $\Sigma^{*}$ is the true covariance matrix of data. Note that the rotation method proposed in Jennrich and Bentler, (2012) completely fails in this case. Before giving the statement of the theorem, we introduce some additional notations. For any matrix $A=(a_{i,j})_{i=1,\ldots,m,j=1,\ldots,n}$, $\mathcal{S}_{1}\subset\{1,\ldots,m\}$ and $\mathcal{S}_{2}\subset\{1,\ldots,n\}$, let $A[\mathcal{S}_{1},\mathcal{S}_{2}]=(a_{i,j})_{i\in\mathcal{S}_{1},j\in\mathcal{S}_{2}}$ be the submatrix of $A$ consisting of elements that lie in rows belonging to set $\mathcal{S}_{1}$ and columns belonging to set $\mathcal{S}_{2}$. For example, consider a matrix $A$ with more than two rows and three columns. With index sets $\mathcal{S}_{1}=\{1,2\}$ and $\mathcal{S}_{2}=\{1,3\}$, the submatrix $A[\mathcal{S}_{1},\mathcal{S}_{2}]$ is a two-by-two matrix, taking the form
$A[\mathcal{S}_{1},\mathcal{S}_{2}]=A[\{1,2\},\{1,3\}]=\left(\begin{array}[]{cc}a_{11}&a_{13}\\ a_{21}&a_{23}\end{array}\right).$
For any set $\mathcal{S}$, let $|\mathcal{S}|$ be the cardinality of $\mathcal{S}$.
@@
Let $\Lambda^{*}$, $\Phi^{*}$ and $\Psi^{*}$ be the true values of the corresponding parameter matrices. The following conditions are sufficient for the identifiability of the bi-factor structure and its parameters.
@@
Given $S=\Sigma^{*}=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$. Suppose that there exists another pair of parameters $\Lambda,\Phi,\Psi$ satisfy the bi-factor structure constraints, we have $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}=\Lambda\Phi(\Lambda)^{\top}$ and $\Psi^{*}=\Psi$.
@@
$|\mathcal{H}^{*}|\geq 2$. In addition, there exists $g_{1}\in\mathcal{H}^{*}$ such that $|\mathcal{B}_{g_{1}}^{*}|\geq 3$ and any 2 rows of $\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}]$ are linearly independent.
@@
Suppose that Conditions 1 and 2 hold. For any parameters $\Lambda,\Phi,\Psi$ that satisfy $S=\Sigma^{*}=\Lambda\Phi(\Lambda)^{\top}+\Psi$, there exist a diagonal sign-flip matrix $D\in\mathcal{D}$ and a permutation matrix $P\in\mathcal{P}$ such that $\Lambda=\Lambda^{*}PD$ and $\Phi=DP^{\top}\Phi^{*}PD$.
The proof of Theorem 1 is given in Appendix G.1.
@@
We note that without additional information, the best we can achieve is to recover $\Lambda^{*}$ and $\Phi^{*}$ up to $\Lambda=\Lambda^{*}PD$ and $\Phi=DP^{\top}\Phi^{*}PD$, where the permutation matrix $P$ and sign-flip matrix $D$ are necessary to account for the label and sign indeterminacies of the factor model. In that case, $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}=\Lambda\Phi(\Lambda)^{\top}$, and thus, the model implied covariance matrix is the same. A similar indeterminacy issue also appears in exploratory factor analysis; see, e.g., Remark 1 in Liu et al., (2023).
@@
Condition 1 ensures the separation between low rank matrix $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$ and diagonal matrix $\Psi^{*}$. A sufficient condition for Condition 1 that can be easily checked in practice is given in Condition 3 below, which requires that each group has at least three non-zero group loadings and there exist at least three groups whose group loadings are linearly independent of the corresponding common loadings. We refer to Theorem 5.1 in Anderson and Rubin, (1956) and Theorem 2 in Fang et al., (2021) for alternative sufficient conditions of Condition 1.
@@
$|\mathcal{B}^{*}_{g}|\geq 3$ for all $g=1,\ldots,G$ and $|\mathcal{H}^{*}|\geq 3$.
@@
Condition 2 is similar to Condition E3S for Proposition 1 in Fang et al., (2021), where the latter is used to ensure the identifiability of parameters when the bi-factor structure is known. It is a sufficient condition that ensures if there is a pair of $\Lambda$ and $\Phi$ that also satisfies the constraints of a bi-factor model and $\Lambda\Phi\Lambda^{\top}=\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$, then there must be a permutation matrix $P$ and sign-flip matrix $D$ such that $\Lambda=\Lambda^{*}PD$ and $\Phi=DP^{\top}\Phi^{*}PD$. This condition first requires the existence of at least two group factors, for each of which the group loadings are linearly independent of the corresponding common loadings. It further requires that there exists a group factor $g_{1}$ among these group factors, such that (1) $g_{1}$ has at least three nonzero group loadings, and (2) any two-by-two submatrix of $\Lambda^{*}$, whose rows correspond to any two variables loading on $g_{1}$ and columns correspond to the common factor and the group factor $g_{1}$, is of full rank. We note that the requirement of Condition 2 is very mild. In fact, the set of parameters not satisfying this condition has zero Lebesgue measure in the parameter space for bi-factor models satisfying that there are at least two group factors $g_{1}$ and $g_{2}$ such that $|\mathcal{B}^{*}_{g_{1}}|\geq 3$ and $|\mathcal{B}^{*}_{g_{2}}|\geq 2$.